OpenStudy (anonymous):

mila plants 820 hectares of corn and potatoes. to maximise his profits he plants 120 hectares more of potatoes than corn. how much of each crop did he plant?

OpenStudy (anonymous):

Let c be the number of corn he plants. potatoes=120+c 120+c+c=820 120+2c=820 Find c. After you've found c, add 120 and you'll get the number of potatoes he planted. Does this make sense?

OpenStudy (anonymous):

yes thank you :)

OpenStudy (anonymous):

you're welcome.